1/3(r + 16) = 1/6 (r + 8)
Answer by checkley71(8403)

1/3(R+16)=1/6(R+8) OR 1/(3R+48)=1/(6R+48) OR 3R+48=6R+48 OR 0=3R OR R=0
PROOF 1/3(0+16)=1/6(0+8) OR 1/48=1/48